2. Ancestral Names Given a list of strings comprised of a name and a Roman numeral, sort the list first by name, then by deciConstraints • 1 sn s 50 • Each names[i] is a single string composed of 2 space-separated values: givenName and romanNumeral.Sample Case o Sample Input STDIN Function 2 → names [] size n = 2 Louis IX → names = [Louis IX, Louis VIII] Louis VIII Sa

Can someone code this asap? Use any language that you want.
Answer #1

As I can use any language, I've used Python

********************

def toDecimal(s): #method to convert roman to decimal
ones={"I":1,"II":2,"III":3,"IV":4,"V":5,"VI":6,"VII":7,"VIII":8,"IX":9} #list storing decimal values for 1-9
tot=0 #value in decimal
pos=0 #position variable
if(s.startswith("L")): #if roman is L i.e., 50
tot=50
return tot #return as 50 is maximum
elif(s.startswith("XL")): #if its 40
tot=40
pos=2
elif(s.startswith("XXX")):
tot=30
pos=3
elif(s.startswith("XX")):
tot=20
pos=2
elif(s.startswith("X")):
tot=10
pos=1
if(pos>=len(s)): #if entire roman string is traversed, return value
return tot
tot=ones[s[pos:]] + tot #else find value for remaining part and add it
return tot

def sortRoman(names):#method to sort name and roman
l=len(names)
nam=[] #three lists for storing name,roman, and decimal
rom=[]
dec=[]
for i in range(l):
det=names[i].split(" ") #split input into name and roman
nam.append(det[0])
rom.append(det[1])
dec.append(toDecimal(rom[i])) #add decimal of roman in list
namdec=[] #new list which consists of names appended with decimal
for i in range(l):
namdec.append(nam[i] + str(dec[i])) #appending name and decimal
(nd,n,r,d)=list(zip(*sorted(zip(namdec,nam,rom,dec)))) #sorting all the lists according to the namedec list
nam=list(n) #converting them to list
rom=list(r)
dec=list(d)
returnlist=[] #new list which will store output
for i in range(l):
returnlist.append(nam[i] + " " + rom[i]) #as the lists are sorted, just append name with roman (as in input)
return returnlist

Code screenshot for indentation help:

def toDecimal(3): #method to convert roman to decimal ones={I:1,II:2,III:3,IV:4,V:5,VI:6,VII:7,VIII:8,IX:9}

def sortRoman (names) : #method to sort name and roman l=len (names) nam=[] #three lists for storing name, roman, and decimal

Driver code:

k=[Steven XL, Steven XVI, David IX, Mary XV, Mary XIII, Mary XX] print (Original list:\n, k) print(\nSorted li

Sample execution:

Original list: [Steven XL, Steven XVI, David IX, Mary XV, Mary XIII, Mary XX] Sorted list: [David IX, Mary XII

In case of any doubt, drop a comment and I'll surely get back to you.

Please give a like if you're satisied with the answer. Thank you.